Allison:

2x - 19, when x = -11.

Allison:

@Vocaloid

Allison:

-41?

Vocaloid:

yes good

Allison:

Ooo good okay.

Allison:

One more, if that's okay?

Vocaloid:

sure

Allison:

p - 3q when p = 9.3 and q = -2.1 I got 15.6 o: Positive~

Vocaloid:

yes good

Allison:

A8Q0WI66SDKphDkO-c4iMQ.png

Allison:

Would this be -8/4 which would be -2?

Vocaloid:

-18-10 = ?

Allison:

-8

Vocaloid:

try factoring out the negative sign

Vocaloid:

-(18+10) = ?

Allison:

8?

Allison:

- 28?!

Vocaloid:

good, now keep going

Allison:

James hush.

Allison:

-28/4 is a negative seven or positive?

Vocaloid:

negative number divided by positive = negative

Vocaloid:

so -7

Allison:

Find the mean of the set of numbers. -9, -8, 4, -1 3.5? Or -3.5? o.o

Vocaloid:

what did you get as your sum?

Allison:

-14?

Vocaloid:

good, now divide that by 4

Vocaloid:

remember our rule for signs

Allison:

-3.5 c:

Vocaloid:

good, yes

Vocaloid:

that's the answer

Allison:

-9qv_odcQ5aVKTWh1mLuHA.png B?

Vocaloid:

we're dividing by negative 1

Vocaloid:

so which one do you think it might be? judging by the names?

Allison:

D

Vocaloid:

"dividing by negative 1"

Allison:

A lololol I just read it~

Vocaloid:

yes good

Allison:

TnC2Ei63Sb698ojoJsv7Mw.png

Allison:

I dunno dis D:

Vocaloid:

we can use process of elimination

Vocaloid:

we're not multiplying by 0, so let's eliminate b

Vocaloid:

there are no reciprocals involved, so eliminate d

Allison:

B, D, gone~~

Allison:

A? .-.

Vocaloid:

we're not multiplying by 1 either, so the commutative property is the best choice

Vocaloid:

tbh this isn't a very good question

Allison:

It confused me D:

Allison:

tABxy_5aTKSoR5bvwWfUfQ.png

Allison:

I chose B~

Vocaloid:

yes good

Owen:

Correct, I just did the math

Allison:

ZOfur5npS0qOq94fB_Q44A.png I chose D because it looked right.

Vocaloid:

yeah that's right

Vocaloid:

the key is that 2/3 and 3/2 cancel out

Vocaloid:

and the two negative signs cancel as well

Allison:

(-16.1) * 9.4 * (-7.2) = x * (-7.2) * (-16.1) A. 9.4 B. 1 C. -16.1 D. -7.2 A?

Vocaloid:

yup good

Allison:

I feel freaking smart.

Shadow:

Because you are

Allison:

Awe c:

Allison:

-9b = -45 B = 5? or -5 x.x

Vocaloid:

when you divide both sides by -9 you get (-45/-9) and negative/negative = positive

Vocaloid:

so 5

Allison:

I always get lost with that.

Allison:

y/6 = -2 -12? :D

Vocaloid:

yes good

Allison:

n/-2 = 24 n = -47?

Allison:

Oops, -48**

Vocaloid:

yes good

Allison:

n/-3 = -0.9 n = -2.7?

Vocaloid:

negative * negative = positive

Allison:

Gahhh.

Vocaloid:

so positive 2.7

Allison:

Got it

Allison:

You're not busy right?

Vocaloid:

nah I'm good

Allison:

Okay

Allison:

xqKYPtZ1ROKzIJ049RH-Dg.png

Allison:

Oh can you see it o.o

Vocaloid:

yeah that's correct

Vocaloid:

yes good

Vocaloid:

yes good

Vocaloid:

you are given d = 41 and t = 2.3 solve for r given d = rt

Allison:

What is R doe?

Vocaloid:

rate "r" is your unknown, and you need to find what it is equal to

Allison:

OH IS IT B?!!? 17.83?!

Vocaloid:

yes good

Vocaloid:

you are given c = g/n where c = 104.5 and n = 6 find g

Allison:

627?

Vocaloid:

yes good

Allison:

http://prntscr.com/gw9o7x p=mv o.0

Vocaloid:

try solving the equation for m

Vocaloid:

how do we get "m" by itself on one side?

Vocaloid:

are there any other instructions?

Vocaloid:

because I see two possible answers

Allison:

I dunno

Allison:

http://prntscr.com/gw9pav

Vocaloid:

oh, in that case, try to get v by itself on one side

Allison:

C? :/

Vocaloid:

p = mv since v is being multiplied by m, let's do the opposite of that. what's the opposite of "multiplying by m"?

Allison:

Dividing by p?

Vocaloid:

almost

Allison:

B?

Allison:

Omg. Okay. D?

Vocaloid:

let's not look at the answer choices for a moment

Allison:

This is making my brain hurt lol

Vocaloid:

if you were typing something in a calculator, and you accidentally multiplied by m, how would you get back to where you were before?

Allison:

Divide by m.

Vocaloid:

good

Allison:

D?

Vocaloid:

so "multiplying by m" is the opposite of dividing by m

Vocaloid:

making D the answer, good

Allison:

http://prntscr.com/gw9qqj Gosh. :/

Vocaloid:

what's the opposite of "dividing by s"?

Allison:

Multiplying by x.

Vocaloid:

almost

Vocaloid:

think back to our calculator example

Vocaloid:

how would you "undo" a division by s?

Allison:

B?

Vocaloid:

think about my question please

Vocaloid:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid how would you "undo" a division by s? \(\color{#0cbb34}{\text{End of Quote}}\)

Allison:

Multiply by s

Vocaloid:

good, so let's take the equation and multiply both sides by s

Vocaloid:

z = x/s zs = x

Vocaloid:

which answer matches?

Allison:

C?

Vocaloid:

x is equal to z multiplied with s

Vocaloid:

try one more time

Allison:

B

Vocaloid:

one more time

Allison:

Oh my God.

Vocaloid:

look for the one where s and z are multiplied together

Allison:

A?

Vocaloid:

yes, good

Allison:

http://prntscr.com/gw9sbw This is getting irritating.

Vocaloid:

what's the opposite of multiplying by mg?

Allison:

Dividing by mg.

Vocaloid:

good, now we divide both sides by mg to get P/mg = h

Vocaloid:

which answer choice matches?

Allison:

C? o.0

Vocaloid:

good

Vocaloid:

can we open a new question, this one is starting to lag

Allison:

Of course.